Atomic weight; Berat atom; Fe =56, Cl= 35.5, 0 = 16, H=l, S=32 C=l2, Na=23, Ca=40, Si=28, P=3 l]g/mol 

Identify the molar mass.

FeCb

H2SO4


1
Expert's answer
2021-11-24T10:24:02-0500

FeCb=56+12+10.811=78.811g

H2SO4=2+32+64=98g